The tax liability for a single individual with taxable income of $67,293 is: options: $7,685. $10,663. $9,239. $10,659?
Answer #1

For single individual tax filer :

$0 - $9700 ------------------------ 10 %

$ 9701- $ 39475 -------------------12%

$ 39,476 - $ 84200 -----------------22%

The above individual has $ 67,293 as the taxable income .

(Slab rate taxation system)

First 9700 taxed @ 10% = 9700 *10%   970

9701 to 39,475 taxed @ 12% = (39475 - 9701 ) * 12%

= 29774 * 12% 3572.88

39476 to 84200 taxed at 22% =( 67293 - 39476 ) * 22%   

=27817 * 22 % 6119.74

Total tax ( 970+3572.88 + 6119.74 ) = 10,662.62

approx. $ 10,663

Thus Answer =$ 10663

If you are benefited from the solution then please LIKE , if disliked then please specify the reason.